OpenStudy (anonymous):

Calculus: Integrals, antiderivatives, and Solving Differential Equations Symbolically:Separation of variables. y'=2xsecy solve for y

OpenStudy (earthcitizen):

http://www.wolframalpha.com/input/?i=y%27%3D2xsecy

OpenStudy (anonymous):

\[\frac{dy}{sec{y}}=2x dx\] \[\int cos(y)dy=\int 2xdx\] \[sin {y} = x^2 +c\] \[y=sin^{-1}{x^2+c}\]

OpenStudy (anonymous):

wait how do you get from step one to step two?

OpenStudy (anonymous):

\[\frac{1}{sec{y}}=cos{y}\] It's a trig identity http://en.wikipedia.org/wiki/Trigonometric_functions

OpenStudy (anonymous):

oh! okay thank you so much :))